# Regularization implies balancedness in the Deep Linear Network

### v1

@@ Abstract.
We use geometric invariant theory (GIT) to study the deep linear network (DLN). The Kempf-Ness theorem is used to establish that the $L^{2}$ regularizer is minimized on the balanced manifold. This allows us to decompose the training dynamics into two distinct gradient flows: a regularizing flow on fibers and a learning flow on the balanced manifold. We show that the regularizing flow is exactly solvable using the moment map.
This approach provides a common mathematical framework for balancedness in deep learning and linear systems theory. We use this framework to interpret balancedness in terms of model reduction and Bayesian principles.
@@ 1. Overview
1.1. The main result
This paper is the second of a series on the mathematical structure of the Deep Linear Network (DLN). We refer to [28] for an introduction and further context.
We study a minimum principle for balancedness that reveals a ‘hidden convexity’ in deep learning. This result contrasts two different geometric structures: the fibers and the balanced varieties. The fiber $\mathcal{F}_{X}$ over an end-to-end matrix $X$ is the algebraic variety defined by the polynomial equation
@@ 1.1. The main result
The balanced variety $\mathcal{M}_{0}$ consists of matrices $\mathbf{W}=(W_{N},\ldots,W_{1})\in\mathbb{M}_{d}^{N}$ such that
$W_{k+1}^{*}W_{k+1}=W_{k}W_{k}^{*},\quad 1\leq k\leq N-1.$ (1.2)
We use to denote the conjugate transpose so that we may study matrices with real and complex entries together. The balanced variety is foliated by rank into a collection of manifolds. When $X$ has full rank, it lies on a leaf of $\mathcal{M}_{0}$, termed the balanced manifold $\mathcal{M}$. The fibers and balanced manifold are illustrated schematically in Figure 1.1. We assume throughout this paper that $X$ has full rank in order to illustrate the new ideas without technical complications.
By hidden convexity we mean that the balanced manifold can be characterized by a class of minimum principles of which the following is the simplest. Consider the $L^{2}$ (ridge) regularizer
@@ 1.2. Balancedness, regularization and Occam’s razor
Similarly, $\mathcal{F}_{X}\cap\mathcal{M}:=\mathcal{O}_{X}$ is a $U_{d}^{N-1}$ group orbit, where $U_{d}$ is the unitary group. Each $\mathbf{W}\in\mathcal{O}_{X}$ is obtained by the group action $\mathbf{Q}\cdot\mathbf{C}$ where $\mathbf{Q}=(Q_{N-1},\ldots,Q_{1})\in U_{d}^{N-1}$ (this is an easy modification of [28, §4]).
@@ 1.3. Balancedness in deep learning and linear systems theory
The concept of balancedness has arisen independently in linear systems theory and deep learning. Our main insight is that these concepts may be unified, allowing us to transport techniques used in linear systems theory to the DLN. In particular, we follow the work of Helmke to prove Theorem 1 [14].
@@ 1.3.1. Linear systems theory
Model reduction in this context is the choice of an optimal realization consistent with the data. Norm balanced realizations minimize the Frobenius norm $\|MAM^{-1}\|_{2}^{2}+\|MB\|_{2}^{2}+\|CM^{-1}\|_{2}^{2}$ over $M\in GL(n;\mathbb{C})$. They constitute a principled choice of an optimal realization and have several favorable properties [14].
@@
The concept of balancedness for the DLN was introduced by Arora, Cohen and Hazan in [1]. The underlying heuristic that ‘load is equally distributed across a balanced network’ was formalized by Du, Hu and Lee for fully connected feed-forward networks with a homogeneous nonlinearity [11, Theorem 2.1]. In our notation, this is the observation that when $\mathbf{W}\in\mathcal{M}$, then $\|W_{k}\|_{2}^{2}$ is independent of $k$. However, for the DLN, more is true. The singular values and singular vectors are aligned across the network: the SVD of $W_{k}=U_{k}\Lambda_{k}V_{k}^{*}$ and $W_{k+1}$ are related through $\Lambda_{k}=\Sigma^{1/N}$ for all $k$ and $V_{k}=U_{k+1}$ for $1\leq k\leq N-1$. This notion of alignment was examined by Ji and Telgarsky in several instances [17]. The relationship between balancing and regularization appears also in the work of Soltanolkotabi, Stöger, and Xie (for $N=2$ and $W_{2}=W_{1}^{*})$ [35].
The surprising appearance of the conservation laws for the DLN (the moments $\mathbf{G}$ defined in equation (1.10) below) has also attracted attention. In several recent papers, Marcotte, Gribonval and Peyré have studied the relation between the symmetries and conservation laws for various neural networks [23, 24, 25]. Minimum principles for balancing weights, including an algorithm for balancing, have been introduced by Saul [34]. This work draws connections between symmetry in deep learning and mathematical physics in a manner that is similar in spirit to our work. Several other recent works have investigated equivariance and symmetry in deep learning [22, 36, 37]. Finally, we note that the interplay between minimum principles and flatness has been studied by Ding, Drusvyatskiy, Fazel and Harchaoui [10].
Our work does not rely on the techniques in the above papers. However, it builds on these themes. Our main contribution is to provide a rigorous variational principle based on well-founded geometric principles that characterizes the relationship between regularization and learning in the DLN. We see the DLN as a benchmark model that provides insight into the harder challenges of nonlinear networks.
@@ 1.3.3. Summary
In terms of mathematical structure, the variational formulation of balancedness in linear systems theory and the DLN reduces to the minimization of a unitarily invariant squared norm on a group orbit. This problem has been solved by the Kempf-Ness theorem in Geometric Invariant Theory (GIT) [20]. Thus, balancedness theorems in both fields are consequences of the Kempf-Ness theorem.
@@
1.4. GIT, the moment map and duality
The main goal in GIT is to classify the orbit space of a group acting on a vector space [30]. We do not study the orbit space of the DLN in generality in this paper since our goal here is to communicate the main insights regarding the dynamics of regularization in its simplest form. However, we note some important consequences of the general theory.
Define the N1N-1 Hermitian matrices
$G_{k}=W_{k}W_{k}^{*}-W_{k+1}^{*}W_{k},\quad 1\leq k\leq N-1.$ (1.10)
The Hermitian matrices $\{G_{k}\}_{k=1}^{N}$ are conserved under the gradient flow of an arbitrary loss function $E$ [28, Theorem 2]. The appearance of such a large number of conservation laws for a gradient flow is surprising at first sight. Indeed, we expect conservation laws for Hamiltonian systems, not gradient flows!
The Kempf-Ness theorem explains this phenomena. The $\{G_{k}\}_{k=1}^{N-1}$ are obtained from the moment map corresponding to the invariance of $\|\mathbf{W}\|_{2}^{2}$ under the unitary $U_{d}^{N-1}$ action
$\mathbb{M}_{d}(\mathbb{C})^{N}\to\mathrm{Her}_{d}^{N-1},\quad\mathbf{W}\mapsto 2\mathbf{G}:=(G_{N-1},\cdots,G_{1}).$ (1.12)
We explain how these arise in Section 3. It is a fundamental result in symplectic geometry that the image of the $U_{d}^{N-1}$ orbit $\mathbf{U}\cdot\mathbf{W}$ under the moment map is a convex set in $\mathrm{Her}_{d}^{N-1}$ [2]. Thus, the moments $\mathbf{G}\in\mathrm{Her}_{d}^{N-1}$ may be seen as the analogues in deep learning of dual variables in the study of conic programs. The factor of $2$ in (1.12) arises because we use the normalization convention in [31] for the moment map.
1.5. The regularizing flow and the learning flow
Recall that the state space for a gradient flow is a Riemannian manifold. Both $\mathcal{F}_{X}$ and $\mathcal{M}$ are smooth embedded submanifolds of $\mathbb{M}_{d}$ under natural assumptions on $X$ (for example, when $X$ has full rank). Thus, they inherit the Euclidean metric from $\mathbb{M}_{d}$. We denote the resulting Riemannian manifolds $(\mathcal{F}_{X},\iota)$ and $(\mathcal{M},\iota)$ respectively. We study two complementary gradient flows on these manifolds:
@@ 1.5. The regularizing flow and the learning flow
    The regularizing flow: This is the gradient flow of $\|\mathbf{W}\|_{2}^{2}$ on the Riemannian manifold $(\mathcal{F}_{X},\iota)$. We write this flow in the form
    $\dot{\mathbf{W}}=-\mathrm{grad}\,\|\mathbf{W}\|_{2}^{2},\quad\mathbf{W}\in(\mathcal{F}_{X},\iota).$ (1.13)
@@
We show that the regularizing flow is exactly solvable in the following sense.
@@
$\mathbf{G}(t)=\mathbf{G}_{0}e^{-4t},\quad t\geq 0.$ (1.15)
@@
Our regularizing flow (1.13) differs from what has been used in linear systems theory. Gradient flows on $GL(n)$ that balance a triple $(A,B,C)$ (‘balancing flows’) have been studied by Helmke and Moore [15]. These gradient flows were inspired by Brockett’s double-bracket flow on $O_{n}$  [7]. However, these works use the normal metric on $GL(n)$ and $O_{n}$ respectively, not the induced metric $\iota$. In our view, it is necessary to use the induced metric instead because (i) this conforms to the Euclidean metric used in practice for deep learning; (ii) it allows us to include noise in a geometrically natural manner using Riemannian Langevin equations (cf. §1.6.2).
@@
Ness studied the gradient flow
$\dot{\mathbf{W}}=-\mathrm{grad}\,\|\mathbf{G}\|_{2}^{2},\quad\mathbf{W}\in(\mathcal{F}_{X},\iota),$ (1.17)
in her analysis of the relationship between GIT and symplectic geometry. She showed that (the projectivized form of) this flow is a Hamiltonian system [31, §3-7]. It is clear that the functional $\|\mathbf{G}\|_{2}^{2}$ is minimized on the balanced manifold. Nevertheless, this flow is of intrinsic mathematical interest and provides a technical relationship between deep learning and mathematical physics through the common structure of Yang-Mills theory [3]. We note that mathematical physics techniques such as the renormalization group and diagrammatic expansions have been used to study deep learning, but rigorous mathematical justification of these ideas is still limited [33]. For these reasons, we present a description of this flow in coordinates in Section 2.
@@
1.6. Summary: a new dynamic paradigm
1.6.1. Gradient flow of a regularized cost function
This is the gradient flow on $\mathbb{M}_{d}^{N}$
where the parameter $\kappa>0$ controls the strength of the regularization. The main observation then is that this dynamical system may be naturally decomposed at each point $\mathbf{W}$ into two orthogonal flows, one normal to $\mathcal{F}_{X}$ (learning) and the other parallel to $\mathcal{F}_{X}$ (regularizing). Our heuristic idea is that regularization is ‘fast’ because of the exponential rate of convergence provided by Theorem 2 (note that the rate is now $\kappa$, not $4$), so that the dynamics of equation (1.18) may be rigorously approximated by the learning and regularizing flows.
1.6.2. Regularization by background noise
1.6.3. Is deep learning ‘secretly Bayesian’?
1.6.4. Conclusion
We now turn to the proofs of Theorem 1 and Theorem 2.
2. The regularizing flow
In this section, we develop the Riemannian geometry of $\mathcal{F}_{X}$, compute the gradient $\mathrm{grad}\|\mathbf{W}\|^{2}$ and prove Theorem 2. We emphasize concrete matrix computations. In the next section, we place these computations within the abstract conception of the Kempf-Ness theorem to establish Theorem 1.

### v2

@@ Abstract.
We use geometric invariant theory (GIT) to study the deep linear network (DLN). The Kempf-Ness theorem is used to establish that the $L^{2}$ regularizer is minimized on the balanced manifold. We introduce related balancing flows using the Riemannian geometry of fibers. The balancing flow defined by the $L^{2}$ regularizer is shown to converge to the balanced manifold at a uniform exponential rate. The balancing flow defined by the squared moment map is computed explicitly and shown to converge globally.
This framework allows us to decompose the training dynamics into two distinct gradient flows: a regularizing flow on fibers and a learning flow on the balanced manifold. It also provides a common mathematical framework for balancedness in deep learning and linear systems theory. We use this framework to interpret balancedness in terms of fast-slow systems, model reduction and Bayesian principles.
@@ 1. Overview
1.1. The main results
This paper is the second of a series on the mathematical structure of the Deep Linear Network (DLN). Our goal in these works is to shed new light on the training dynamics on deep learning through an analysis of the hidden symmetries of the DLN. This paper focuses on the concept of balancedness; we observe the natural role of Geometric Invariant Theory (GIT) in the DLN and use the Kempf-Ness theorems and related gradient flows to establish a new dynamic paradigm in deep learning.
We summarize the essentials of the DLN below and refer to [35] for further context. The state space of the DLN consists of NN $d\times d$ matrices $\mathbf{W}=(W_{N},\ldots,W_{1})\in\mathbb{M}_{d}^{N}(\mathbb{F})$. The integers $N$ and $d$ are termed the depth and width respectively. The field $\mathbb{F}$ may be either $\mathbb{R}$ or $\mathbb{C}$ and we use to denote the transpose or conjugate transpose respectively. The cost function $E$ for the DLN depends only on the end-to-end matrix
@@ 1.1. The main results
Training dynamics in the DLN is modeled by the gradient flow
$\dot{\mathbf{W}}=-\nabla_{\mathbf{W}}E(X(\mathbf{W})).$ (1.2)
The underlying Euclidean metric is described in Section 1.5 below.
The natural geometry of equation (1.2) is that of a gradient flow on a fibered space. It is this aspect that we study in detail. Our results in this paper rely on the interplay between two complementary foliations of $\mathbb{M}_{d}^{N}$:
  1. (1)
    The fibers $\mathcal{F}_{X}$ of $\mathbb{M}_{d}^{N}$ defined as the solution set to the polynomial equation (1.1) for each $X\in\mathbb{M}_{d}$.
  2. (2)
    the balanced variety $\mathcal{M}_{0}$ defined as the solution set to the quadratic system
    $W_{k}W_{k}^{*}=W_{k+1}^{*}W_{k+1},\quad 1\leq k\leq N-1.$ (1.3)
The balanced variety is invariant under the gradient flow (1.2) and has several fundamental properties (see [35, Thms. 2-8])). It is foliated by rank into a collection of manifolds. When $X$ has full rank, it lies on a leaf of $\mathcal{M}_{0}$, termed the balanced manifold $\mathcal{M}$. We assume throughout this paper that $X$ has full rank in order to illustrate the new ideas without technical complications. The fibers and balanced manifold are illustrated schematically in Figure 1.1.
We establish a minimum principle for balancedness that reveals a ‘hidden convexity’ in deep learning. By hidden convexity we mean that the balanced manifold can be characterized by a class of minimum principles of which the following is the simplest. Consider the $L^{2}$ (ridge) regularizer
@@
In what follows, we provide several interpretations and extensions of this theorem. The most important in practice is model reduction: Theorem 1 selects the simplest parametric description of a given end-to-end matrix $X$ (Section 1.2). Our ideas on balancing are closely tied to linear systems theory; we show that both areas may be unified through the use of GIT (Sections 1.31.4). In order to implement this idea in practice, we introduce a broader class of minimization principles along with new gradient flows (regularizing flows) with rigorous convergence guarantees (Section 1.5). Finally, we use GIT to relate the DLN to mathematical physics and the Yang-Mills theory. Let us now explain these ideas in greater detail.
@@ 1.2. Balancedness, regularization and Occam’s razor
Similarly, $\mathcal{F}_{X}\cap\mathcal{M}:=\mathcal{O}_{X}$ is a $U_{d}^{N-1}$ group orbit, where $U_{d}$ is the unitary group. Each $\mathbf{W}\in\mathcal{O}_{X}$ is obtained by the group action $\mathbf{Q}\cdot\mathbf{C}$ where $\mathbf{Q}=(Q_{N-1},\ldots,Q_{1})\in U_{d}^{N-1}$ (this is an easy modification of [35, §4]).
@@ 1.3. Balancedness in deep learning and linear systems theory
The concept of balancedness has arisen independently in linear systems theory and deep learning. Our main insight is that these concepts may be unified, allowing us to transport techniques used in linear systems theory to the DLN. In particular, we follow the work of Helmke to prove Theorem 1 [18]. We augment his work with an analysis of the Riemannian geometry of $\mathcal{F}_{X}$ and a rigorous analysis of balancing flows, providing an independent proof of Theorem 1.
@@ 1.3.1. Linear systems theory
Model reduction in this context is the choice of an optimal realization consistent with the data. Norm balanced realizations minimize the Frobenius norm $\|MAM^{-1}\|_{2}^{2}+\|MB\|_{2}^{2}+\|CM^{-1}\|_{2}^{2}$ over $M\in GL(n;\mathbb{C})$. They constitute a principled choice of an optimal realization and have several favorable properties [18]. This concept continues to be of great importance in Structured State Space Model (SSM) architectures, which are principled alternative to transformer architectures for modeling long sequences [17]. In particular, the analysis of implicit bias in SSMs in analogy with implicit bias in the DLN is of current importance [12, 39].
@@
The concept of balancedness for the DLN was introduced by Arora, Cohen and Hazan in [2]. The underlying heuristic that ‘load is equally distributed across a balanced network’ was formalized by Du, Hu and Lee for fully connected feed-forward networks with a homogeneous nonlinearity [14, Theorem 2.1]. In our notation, this is the observation that when $\mathbf{W}\in\mathcal{M}$, then $\|W_{k}\|_{2}^{2}$ is independent of $k$. However, for the DLN, more is true. The singular values and singular vectors are aligned across the network: the SVD of $W_{k}=U_{k}\Lambda_{k}V_{k}^{*}$ and $W_{k+1}$ are related through $\Lambda_{k}=\Sigma^{1/N}$ for all $k$ and $V_{k}=U_{k+1}$ for $1\leq k\leq N-1$. This notion of alignment was examined by Ji and Telgarsky in several instances [21]. The relationship between balancing and regularization appears also in the work of Soltanolkotabi, Stöger, and Xie (for $N=2$ and $W_{2}=W_{1}^{*})$ [44].
The surprising appearance of the conservation laws for the DLN (the moments $\mathbf{G}$ defined in equation (1.11) below) has also attracted attention. In several recent papers, Marcotte, Gribonval and Peyré have studied the relation between the symmetries and conservation laws for various neural networks [30, 31, 32]. Minimum principles for balancing weights, including an algorithm for balancing, have been introduced by Saul [42]. This work draws connections between symmetry in deep learning and mathematical physics in a manner that is similar in spirit to our work. Several other recent works have investigated equivariance and symmetry in deep learning [27, 45, 48]. Finally, we note that the interplay between minimum principles and flatness has been studied by Ding, Drusvyatskiy, Fazel and Harchaoui [13] and that a stratification of the loss landscape for the quadratic loss function has been studied by Achour, Malgouyres and Gerchinovitz [1].
Our work does not rely on the techniques in the above papers though it builds on these themes. Instead, our work is closest in spirit to the relationship between optimization problems and integrable gradient flows on adjoint orbits studied by Bloch, Brockett and collaborators in the 1990s [8, 7]. Recent developments in this vein have extended these geometric principles to high-dimensional machine learning and quantum control, reinforcing the notion that the training dynamics of deep architectures are fundamentally governed by the underlying Riemannian geometry and conservation laws induced by parameter space symmetries [47].
1.4. GIT, the moment map and duality
The main goal in Geometric Invariant Theory (GIT) is to classify the orbit space of a group acting on a vector space [37]. We do not study the orbit space of the DLN in full generality. However, we note some important consequences of the general theory.
1.4.1. The moment map
Define the N1N-1 Hermitian matrices
$G_{k}=W_{k}W_{k}^{*}-W_{k+1}^{*}W_{k+1},\quad 1\leq k\leq N-1.$ (1.11)
The Hermitian matrices $\{G_{k}\}_{k=1}^{N}$ are conserved under the gradient flow of an arbitrary loss function $E$ [35, Theorem 2]. The appearance of such a large number of conservation laws for a gradient flow is surprising at first sight. Indeed, we expect conservation laws for Hamiltonian systems, not gradient flows!
The Kempf-Ness theorem explains this phenomena. The $\{G_{k}\}_{k=1}^{N-1}$ are obtained from the moment map corresponding to the invariance of $\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}$ under the action
$\mathbb{M}_{d}(\mathbb{C})^{N}\to\mathrm{Her}_{d}^{N-1},\quad\mathbf{W}\mapsto\mathbf{G}:=(G_{N-1},\cdots,G_{1}).$ (1.13)
We explain how these arise in Section 3. Loosely speaking, the moments $\mathbf{G}\in\mathrm{Her}_{d}^{N-1}$ may be seen as the analogues in deep learning of dual variables in the study of conic programs.
@@ 1.4.2. Summary
In terms of mathematical structure, the variational formulation of balancedness in linear systems theory and the DLN reduces to the minimization of a unitarily invariant squared norm on a group orbit. This problem has been solved by the Kempf-Ness theorem in GIT [24]. Thus, balancedness theorems in both fields are consequences of the Kempf-Ness theorem. However, the Kempf-Ness theory does not provide quantitative rates of convergence or practical balancing algorithms. To this end, we use the Riemannian geometry of fibers to construct balancing flows with rigorous convergence guarantees (Theorem 2 and Theorem 10).
@@
1.5. Regularizing flows and the learning flow
Recall that the state space for a gradient flow is a Riemannian manifold. Thus, in order to define gradient flows on $\mathcal{F}_{X}$ and $\mathcal{M}$, we must equip them with a Riemannian metric.
It is only the complex case that requires some comment. When $\mathbb{F}=\mathbb{C}$, we equip $\mathbb{M}_{d}(\mathbb{C})$ with the inner-product
$\langle A,B\rangle=\mathrm{Re}\,\mathrm{Tr}\left(A^{*}B\right).$ (1.14)
Writing $A=a+ib$ and $B=c+id$ for the real and imaginary parts, we see that
$\langle A,B\rangle=\mathrm{Tr}\left(a^{T}c+b^{T}d\right),$ (1.15)
so that $\mathbb{M}_{d}(\mathbb{C})$ has the same norm as the real vector space $\mathbb{M}_{d}(\mathbb{R})\times\mathbb{M}_{d}(\mathbb{R})$. Similarly, we define the norm on $\mathbb{M}_{d}^{N}(\mathbb{C})$
$\langle\mathbf{A},\mathbf{B}\rangle=\sum_{k=1}^{N}\mathrm{Re}\,\mathrm{Tr}\left(A_{k}^{*}B\right).$ (1.16)
This inner-product defines $\mathbb{M}_{d}^{N}(\mathbb{C})$ as a Euclidean space. Since the fiber $\mathcal{F}_{X}$ and the balanced variety $\mathcal{M}$ are defined as the solutions to the algebraic equations (1.1) and (1.1), they are embedded Riemannian manifolds of the Euclidean space $\mathbb{M}_{d}^{N}(\mathbb{F})$ under natural assumptions on $X$ (for example, when $X$ has full rank). We denote the resulting Riemannian manifolds $(\mathcal{F}_{X},\iota)$ and $(\mathcal{M},\iota)$ respectively.
We study two complementary gradient flows on these manifolds:
@@ 1.5. Regularizing flows and the learning flow
    Regularizing flows: This is the gradient flow of a regularizer $R(\mathbf{W})$ on the Riemannian manifold $(\mathcal{F}_{X},\iota)$. Theorem 1 corresponds to the regularizer $R(\mathbf{W})=\tfrac{1}{2}\|\mathbf{W}\|_{2}^{2}$ and the gradient flow
    $\dot{\mathbf{W}}=-\frac{1}{2}\mathrm{grad}\,\|\mathbf{W}\|_{2}^{2},\quad\mathbf{W}\in(\mathcal{F}_{X},\iota).$ (1.17)
    We also consider the gradient flow of the squared moment map, $R(\mathbf{W})=\tfrac{1}{2}\|\mathbf{G}\|^{2}$ (see Remark 7 below). Theorem 1 extends to include the Schatten $p$-norms as regularizers for $1<p<\infty$. It does not yet include the important case $p=1$ (see Theorem 20 in Section 3).
@@
We show that the regularizing flow for the $L^{2}$-regularizer is exactly solvable in the following sense.
@@
$\mathbf{G}(t)=\mathbf{G}_{0}e^{-2t},\quad t\geq 0.$ (1.19)
@@
The underlying algebraic geometry, including the foliation by rank, has been analyzed in generality by Shewchuk and Bhattacharya [43] and Pepin Lehalleur and Rimányi [28]. While these results parallel one another, the methods are different: the article [43] emphasizes linear algebra, whereas [28] organizes these calculations using the language of quiver representations. The relationship between these geometric structures and the loss landscape for the quadratic loss function is a topic of current interest.
@@
Our regularizing flow (1.17) differs from what has been used in linear systems theory. Gradient flows on $GL(n)$ that balance a triple $(A,B,C)$ (‘balancing flows’) have been studied by Helmke and Moore [19]. These gradient flows were inspired by Brockett’s double-bracket flow on $O_{n}$  [9]. However, these works use the normal metric on $GL(n)$ and $O_{n}$ respectively, not the induced metric $\iota$. In our view, it is necessary to use the induced metric instead because (i) this conforms to the Euclidean metric used in practice for deep learning; (ii) it allows us to include noise in a geometrically natural manner using Riemannian Langevin equations (cf. §4.2).
@@
Kirwan and Ness studied the Morse theory of the squared moment map in their pioneering work on GIT [25, 38]. Ness studied the related gradient flow [38, §3-7] and analyzed its convergence properties. In our context, this regularizing flow, which we term the Kirwan-Ness flow, is given by  [^1] The reader should note that we use a different normalization from Ness. We find it more convenient to use $\tfrac{1}{2}\|\mathbf{G}\|_{2}^{2}$ rather than $\|\mathbf{G}\|_{2}^{2}$.
$\dot{\mathbf{W}}=-\frac{1}{2}\mathrm{grad}\,\|\mathbf{G}\|_{2}^{2},\quad\mathbf{W}\in(\mathcal{F}_{X},\iota).$ (1.21)
A striking feature of the DLN is that this flow is given explicitly by the cubic system
$\dot{W}_{k}=2(W_{k}G_{k-1}-G_{k}W_{k}),\quad 1\leq k\leq N,$ (1.22)
with the convention $G_{N}=G_{0}=0$. This allows us to adapt the general theory from GIT for the purpose of balancing in deep learning (see Section 2.5 below).
The Kirwan-Ness flow also provides a technical relationship between deep learning and mathematical physics through the common structure of Yang-Mills theory as discussed in Uhlenbeck’s recent work [3, 46]. We note that mathematical physics techniques such as the renormalization group and diagrammatic expansions have been used to study deep learning, but rigorous mathematical justification of these ideas is still limited [41].
@@
1.6. Organization of the paper
We first study the Riemannian geometry of fibers, introduce the regularizing flows and establish their convergence (Theorem 2 and Theorem 10) in Section 2. This is followed by a discussion of the Kempf-Ness theorem, its generalization, the Azad-Loeb theorem, and the application of GIT to the DLN. The paper concludes with a discussion of a new dynamic paradigm suggested by the methods of this paper.
2. Regularizing flows
In this section, we develop the Riemannian geometry of $\mathcal{F}_{X}$, compute the gradient $\mathrm{grad}\|\mathbf{W}\|^{2}$ and prove Theorem 2. We emphasize concrete matrix computations. In the next section, we place these computations within the abstract conception of the Kempf-Ness theorem to establish Theorem 1. All calculations are performed under the assumption that $\mathbb{F}=\mathbb{C}$. The case $\mathbb{F}=\mathbb{R}$ follows naturally.